OpenStudy (anonymous):

find the standard deviation : 3 5 9 7 2 6 8 4 5 7

OpenStudy (displayerror):

Find the sample mean: \[\bar{X} = \frac{3+5+9+7+2+6+8+4+5+7}{10} = 5.6\] Recall the formula for standard variance: \[\sigma ^2 = \frac{1}{n-1} \sum_{i=1}^{10} (X_i - \bar{X})^2\] This formula requires that you subtract each value (3, 5, 9, etc.) from the sample mean, square this difference, sum it up, then divide that sum by n-1, where n = the sample size which, in this case, is 10. An example (for the first two numbers) to help get you started: \[\sigma ^2 = \frac{1}{10-1} (3 - 5.6)^2 + (5 - 5.6)^2 + \ldots\] Remember that you're looking for the sample standard deviation, which is the square root of the sample variance (which you found above).

OpenStudy (anonymous):

thank you so much